John borrowed a certain amount of money from Tebogo at a simple interest rate of 8,7% per year.
After five years John owes Tebogo R10 000,00. Calculate how much money John initially borrowed
1
Expert's answer
2021-07-06T05:26:40-0400

Calculation of the princ ipal loan amount:-


Amount = Principal (1 + Rate X Time).


R1000,000 = P (1 + 0.087 X 5).


R1, 000,000 = P (1 + 0.435).


Principal loan "=\frac {R1000,000} {1.435}"


Principal loan = R696,864.11.